# ILCL: Inverse Logic-Constraint Learning from Temporally Constrained Demonstrations

### v1

@@ IV-B Constraint player: GA-based temporal-logic mining
At each $k$-th iteration of ILCL, the constraint player—GA-TL-Mining—identifies a TLTL constraint $\phi$ that maximizes a fitness score $\mathcal{F}_{k}(\phi)$ given demonstrations $\Xi_{E}$. The score $\mathcal{F}_{k}(\phi)$, a Monte Carlo approximation of the dual objective in Eq. (4), increases with the constraint $\phi$ that penalizes violations by prior policies $\pi_{1:k}$, while assigning zero if violated by expert trajectories $\Xi_{E}$; We define the fitness score $\mathcal{F}_{k}(\phi)$ as
$\frac{\sum_{\xi\in\Xi_{1:k}}\mathds{1}[\rho(\xi,\phi)<0]}{|\Xi_{1:k}|}\cdot\Pi_{\xi_{E}\in\Xi_{E}}\mathds{1}[\rho(\xi_{E},\phi)>0],$ (6)
where $\Xi_{1:k}=\{\xi\,|\,\xi\sim\pi_{i},i\in[1,k]\}$ and $\pi_{i}$ denotes policies generated by Logic-CRL. Using the fitness function, GA-TL-Mining iteratively improves a population of TLTL constraints, represented as logic trees, through selection, crossover, and mutation, following the tree-based GA framework, ROGE [22]. For simplicity, we refer TLTL constraints and logic trees interchangeably.
GA-TL-Mining consists of three steps—1) initial population generation, 2) parent selection, and 3) offspring generation. We describe detail below.
@@
We then construct $N_{\text{R}}$ random logic trees by modifying the basis trees sampled to enhance the diversity of the initial population. For each, we begin by randomly selecting a basis tree and iteratively injecting a randomly chosen operator as the parent of a randomly selected node. This process continues until the tree exceeds a predefined depth $d_{R}$, or terminates early with probability $p_{R}$. In this work, we set $d_{R}=3$ and $p_{R}=0.1$. Finally, the size of the population becomes $N=N_{\text{B}}+N_{\text{R}}$. Note that, when $k>1$, the initial population includes the previously discovered $k-1$ TLTL constraints, replacing $k-1$ random trees to guide subsequent mining.
2) Parent selection: This step selects $N_{p}$ parent logic trees based on the highest regularized fitness scores within the current population $\Phi=\{\phi_{1},\ldots,\phi_{N}\}$ to generate offspring for the next generation. Let $\mathcal{V}(\phi_{i})$ denote the node set of each tree $\phi_{i}$. The regularized fitness score $\mathcal{F}_{k}^{\Phi}(\phi)$ is
@@
where $\zeta$ is a hyperparameter and the subtraction of $2$ accounts for the minimum number of nodes in a tree. To evaluate $\mathcal{F}_{k}^{\Phi}(\phi)$, we first compute the fitness score $\mathcal{F}_{k}$ by optimizing the parameters $\boldsymbol{\theta}$ of $\phi$ using dual annealing , as described in [5]. We then regularize the score by penalizing large tree sizes, encouraging the discovery of simpler constraints in this ill-posed mining problem [22].
Figure 3: The illustration of the operations in the offspring generation. For each pTLTL tree in the figure, the blue nodes represent logical operators, and the orange nodes represent pAP nodes. The labels of the pAP node indicate the semantics of pAP; for example, the label ’2:<’ refers to ${\bf s}^{(2)}<\theta$ for an undetermined parameter $\theta$.
3) Offspring generation: We iteratively generate $N$ offspring of the current population $\Phi=\{\phi_{1},\dots,\phi_{N_{p}}\}$ applying three genetic operators: CROSSOVER, MUTATION-R, and MUTATION-A (see Fig. 3). For each iteration, we first randomly select two parent trees, $\phi_{i}$ and $\phi_{j}$, using tournament selection We then randomly apply an genetic operator to generate a new offspring in the following manner:
@@
  • MUTATION-R: we first randomly select a node in the tree $\phi_{i}$. We then either replace it with another node of the same type—pAP or logic operator—or remove it along with its descendants.
  • MUTATION-A: we insert a randomly selected logic operator as the parent of a randomly chosen node in the tree $\phi_{1}$ preserving semantics. For example, we mutate $\mu_{1}\,\mathcal{U}\,\mu_{2}$ into $\mu_{1}\,\mathcal{U}\,\square\mu_{2}$.
[…]
@@ IV-C Policy player: logic-constrained reinforcement learning
We introduce logic-constrained RL (Logic-CRL), which trains a policy to satisfy the TLTL constraints discovered by GA-TL-Mining. A main challenge is that TLTL constraints are non-Markovian; Their truths depend on entire state sequences rather than individual states. To resolve the mismatch, we embed each constraint into a product CMDP (PCMDP) by augmenting states with the TLTL constraint-based automaton states, thereby restoring the Markov property. Another challenge is the non-differentiable nature of logic constraints. To figure it out, we adopt Lagrangian CRL [25] with a novel state-wise constraint redistribution scheme, enabling conventional CRL frameworks to simplify the non-differentiable problem while still satisfying the non-Markovian TLTL constraint. We detail both processes below.
Product Constrained Markov Decision Process: We construct a product CMDP, we call PCMDP, by synchronously combining the CMDP’s transition system with a deterministic finite automata (DFA) converted from a TLTL constraint $\phi$ The DFA is the tuple $(\mathcal{Q}_{\phi},\Sigma_{\phi},\delta_{\phi},q_{0,\phi},F_{\phi})$, where $\mathcal{Q}_{\phi}$ is the set of states, $\Sigma_{\phi}=2^{\textbf{AP}}$ is the power set of APs, $\delta_{\phi}:\mathcal{Q}_{\phi}\times\Sigma_{\phi}\rightarrow\mathcal{Q}_{\phi}$ is the deterministic transition function, $q_{0,\phi}\in\mathcal{Q}_{\phi}$ is the initial state, and $F_{\phi}\subset\mathcal{Q}_{\phi}$ is the set of accepting states.
@@
Finally, PCMDP enables the synthesis of a policy that takes the TLTL constraint $\phi$ into account while maximizing accumulated rewards. However, optimizing policies with TLTL constraints presents challenges due to their sparse nature and the delayed evaluation at the end of episodes.
Lagrangian-based Constrained Reinforcement Learning: We introduce a Lagrangian CRL resolving the problem of sparse and temporally delayed evaluation of TLTL constraints in policy learning. The sparse nature comes from Boolean semantics of TLTLs. To figure it out, we incorporate quantitative semantics with robustness function $\rho(\xi,\phi)$ to define a bounded but dense constraint-cost function $C_{\phi}^{\prime}\in[0,1]$, updating the discrete function $C_{\phi}$ in Eq. 5, as:
@@
where $\alpha\in[0,1]$ is a mixing weight, $\Upsilon\in\mathbb{R}^{+}$ limits the impact of large constraint violations, and $\mathrm{CLIP}$ truncates the negative robustness value within the range $[0,\Upsilon]$.
To address the delayed evaluation, we introduce a redistribution scheme that spreads out the trajectory-level constraint cost $C^{\prime}_{\phi}(\xi)$ uniformly across all state-action pairs $({\bf s}_{t},{\bf a}_{t})\in\xi$, analogous to the uniform reward redistribution in IRCR [9]. The resulting stepwise constraint cost $C^{\prime}_{\phi}({\bf s}_{t},{\bf q}_{t},{\bf a}_{t})$ is on the augmented state space $({\bf s}_{t},{\bf q}_{t})\in\mathcal{S}_{\phi}$ as
$C^{\prime}_{\phi}({\bf s}_{t},{\bf q}_{t},{\bf a}_{t})=\mathbb{E}_{\xi\sim\mathcal{D}}[C^{\prime}_{\phi}(\xi)|(({\bf s}_{t},{\bf q}_{t}),{\bf a}_{t})\in\xi],\vspace{-5pt}$ (11)
where $\mathcal{D}$ is the replay buffer containing previously collected trajectories in the augmented state space. This trajectory-space smoothing converts the trajectory constraint in Eq. (5) into a state-action constraint, enabling its use in Lagrangian CRL.
We finally formulate the Lagrangian dual problem of Eq. (5) that finds a constrained optimal policy $\pi_{k}$ in the $k$-th Logic-CRL process:
@@
where $\lambda$ is a learnable Lagrangian multiplier and $d$ is a constraint-cost threshold. We define a large threshold, $d=\frac{\varepsilon\alpha}{N_{\xi}}$, for stable optimization while enforcing the satisfaction of the binary logic constraint $C_{\phi}$, where $\varepsilon<1$ is a positive constant and $N_{\xi}$ is the number of trajectories sampled from the current policy $\pi$. To solve Eq. (12), we adopt SAC-Lag [29]—a soft actor-critic (SAC) [32] version of a Lagrangian CRL method.
Further, to enhance off-policy learning in Eq. (12), we initialize the replay buffer $\mathcal{D}$ with the expert trajectories $\Xi_{E}$ and the previously sampled trajectories $\Xi_{1:k-1}$ that already satisfy the given constraint After obtaining $\pi_{k}$, we sample zero-violation trajectories $\Xi_{k}$ and update the trajectory set as $\Xi_{1:k}=\Xi_{1:k-1}\cup\Xi_{k}$, as input to the next GA-TL-Mining step.
Lastly, after $N$ alternations of GA-TL-Mining and Logic-CRL, ILCL returns the best constraint $C_{\phi}^{*}$. Although the two-player game ideally converges, the limited iterations and adversarial instability often yield a suboptimal constraint. Thus, we choose the most expert-like constraint $C_{\phi,i}$ from the obtained constraint set $\{C_{\phi,1},\dots,C_{\phi,N}\}$. For each candidate constraint, we evaluate its associated policy $\pi_{i}$ on the expert trajectories $\Xi_{E}$ and select the policy that minimizes the error to the expert actions,
$\pi_{i}=\mathop{\arg\min}_{\pi_{i}\in[\pi_{1},...,\pi_{N}]}\;\mathbb{E}_{\begin{subarray}{c}(s_{E},a_{E})\sim\Xi_{E}\\ a\sim\pi_{i}(\cdot|s_{E})\end{subarray}}||a_{E}-a||_{2}.$ (13)
We then report $C_{\phi,i}$ corresponding to $\pi_{i}$.
@@ V-A Benchmark Setups
1) Navigation: This task requires a point agent at $\mathbf{x}^{Agt}\in\mathbb{R}^{2}$ navigating to a goal position $\mathbf{x}^{Goal}\in\mathbb{R}^{2}$ while satisfying temporal constraints related to colored regions randomly placed in the workspace (see Fig. 4 (Left)). Let $\mathbf{p}^{R}$, $\mathbf{p}^{G}$, and $\mathbf{p}^{B}$ denote the polar coordinates $(p_{dist},p_{ang})\in\mathbb{R}^{2}$ of the nearest red, green, and blue regions, respectively, measured relative to the agent’s coordinate. For example, ${\bf p}^{R}=(p^{R}_{dist},p^{R}_{ang})$. We model this environment as two CMDPs, each subject to a distinct constraint, $\phi_{\text{GT1}}$ or $\phi_{\text{GT2}}$:
@@
2) Wiping: This task requires a roller agent at ${\bf x}^{Agt}\in\mathbb{R}^{2}$ wiping the ground surface between the start point $A$ (green) and the goal point $B$ (red), as illustrated in Fig. 4 (Middle). Let $f_{c}$ be the contact force between the roller and the surface. ${}^{A}{\bf x}^{Agt}$ and ${}^{B}{\bf x}^{Agt}$ denote the positions of the agent relative to the start and goal points, respectively. For example, ${}^{A}{\bf x}^{Agt}=(^{A}x^{Agt},^{A}y^{Agt})$. We model this environment as a CMDP:
@@
  • $R=R_{goal}+R_{contact}+R_{penalty}$; $R_{goal}=\mathds{1}(|^{B}x^{Agt}|<0.1)-|^{B}x^{Agt}|\cdot\mathds{1}(|^{B}x^{Agt}|>0.1)$ is a goal reward, $R_{contact}=2\cdot\mathds{1}(f_{c}>0.05)$ is a contact reward to encourage wiping contact, and $R_{penalty}=-0.2\cdot f_{c}^{2}$ is to penalize large contact.
@@
3) Peg-in-shallow-hole: This task requires a gripper agent to insert a thin peg into a shallow-angled hole maintaining its side contact as shown in Fig. 4 (Right). Let $\Sigma_{h}$ denote the 2D hole frame, tilted by angle $\theta^{h}$. The task uses a parallel-jaw gripper positioned at ${}^{\Sigma_{h}}{\bf x}^{grip}$, represented by $(^{\Sigma_{h}}x^{grip},^{\Sigma_{h}}y^{grip},^{\Sigma_{h}}\theta^{grip})\in\mathbb{R}^{3}$, where $\theta^{grip}$ represents the tilt angle of the gripper in the hole frame. Likewise, a peg is at ${}^{\Sigma_{h}}{\bf x}^{peg}$. In addition, we introduce two features: $d_{hole}^{peg}$ and $d_{jaw}^{peg}$, which measure the peg edge-to-hole and peg edge-to-support jaw distances, respectively. We then model this environment as a CMDP:
@@
  • $R=(1-\tanh(5|^{\Sigma_{h}}y^{peg}-^{\Sigma_{h}}y^{peg*}|))+5\cdot\mathds{1}(^{\Sigma_{h}}y^{peg}-^{\Sigma_{h}}y^{peg*})$, where ${}^{\Sigma_{h}}y^{peg}$ and ${}^{\Sigma_{h}}y^{peg*}$ are the current and desired vertical position of the peg in terms of the hole frame $\Sigma_{h}$.
@@
Figure 5: Comparison of the proposed ILCL and five baseline methods in training and test environments. ‘Expert’ denotes the result of expert demonstrations without constraint violations. We normalize all REW values to the ‘Expert’ score, resulting in a range of $[0,1]$.

### v2

@@ IV-B Constraint player: GA-based temporal-logic mining
At iteration $k$ of ILCL, the constraint player—GA-TL-Mining—identifies a TLTL constraint $\phi$ that maximizes a fitness score $\mathcal{F}_{k}(\phi)$ given demonstrations $\Xi_{E}$. The score $\mathcal{F}_{k}(\phi)$ is a Monte Carlo approximation of the dual objective in Eq. (4). This increases when $\phi$ penalizes violations by prior policies $\pi_{1:k}$, while assigning zero if violated by expert trajectories $\Xi_{E}$;
$\mathcal{F}_{k}(\phi)\!=\!\frac{\sum_{\xi\in\Xi_{1:k}}\mbox{$\mathbf{1}$}[\rho(\xi,\phi)\!<\!0]}{|\Xi_{1:k}|}\!\cdot\!\Pi_{\xi_{E}\in\Xi_{E}}\mbox{$\mathbf{1}$}[\rho(\xi_{E},\phi)\!>\!0],$ (6)
where $\Xi_{1:k}\!=\!\{\xi|\xi\!\sim\!\pi_{i},i\!\in\![1,k]\}$ and $\pi_{i}$ denotes policies generated by Logic-CRL. Using the fitness, GA-TL-Mining iteratively improves a population of TLTL constraints, represented as logic trees, via selection, crossover, and mutation, following the tree-based GA framework, ROGE [nenzi2018robust]. For simplicity, we refer TLTL constraints and logic trees interchangeably.
GA-TL-Mining consists of three steps—1) initial population generation, 2) parent selection, and 3) offspring generation:
@@
We then construct $N_{\text{R}}$ random logic trees by modifying the basis trees sampled to enhance the diversity of the initial population. For each, we begin by randomly selecting a basis tree and iteratively injecting a randomly chosen operator as the parent of a randomly selected node. This process continues until the tree exceeds a predefined depth $d_{R}$, or terminates early with probability $p_{R}$. In this work, we set $d_{R}=3$ and $p_{R}=0.1$. Finally, the size of the population becomes $N=N_{\text{B}}+N_{\text{R}}$. Note that, when $k>1$, the initial population includes the previously discovered k1k-1 TLTL constraints, replacing k1k-1 random trees to guide subsequent mining.
2) Parent selection: This step selects $N_{p}$ parent logic trees based on the highest regularized fitness scores within the current population $\Phi=\{\phi_{1},\ldots,\phi_{N}\}$ to generate offspring for the next generation. Let $\mathcal{V}(\phi_{i})$ denote the node set of each tree $\phi_{i}$. The regularized fitness score $\mathcal{F}_{k}^{\Phi}(\phi)$ is:
@@
where $\zeta$ is a hyperparameter and the 2-2 term accounts for the minimum number of nodes in a tree. To evaluate $\mathcal{F}_{k}^{\Phi}(\phi)$, we first compute the fitness score $\mathcal{F}_{k}$ by optimizing the parameters $\bm{\theta}$ of $\phi$ via dual annealing [liu2024interpretable]. We then regularize the score by penalizing large tree sizes, encouraging the discovery of simpler constraints in this ill-posed mining problem [nenzi2018robust].
Figure 3: The illustration of the operations in the offspring generation. For each pTLTL tree in the figure, the blue nodes represent logical operators, and the orange nodes represent pAP nodes. The labels of the pAP node indicate the semantics of pAP; for example, the label ’2:¡’ refers to ${\bf s}^{(2)}<\theta$ for an undetermined parameter $\theta$.
3) Offspring generation: We iteratively generate $N$ offspring from the current population $\Phi=\{\phi_{1},\dots,\phi_{N_{p}}\}$ applying three genetic operators: CROSSOVER, MUTATION-R, and MUTATION-A (see Fig. 3). For each iteration, we first randomly select two parent trees, $\phi_{i}$ and $\phi_{j}$, using tournament selection. We then randomly apply a genetic operator to generate a new offspring in the following manner:
@@
  • MUTATION-R: we randomly select a node in $\phi_{i}$ and either replace it with a same-type node—pAP or logic operator—or remove it along with its descendants.
  • MUTATION-A: we insert a randomly selected logic operator as the parent of a randomly chosen node in the tree $\phi_{1}$ preserving semantics. (e.g., $\mu_{1}\,\mathcal{U}\,\mu_{2}$ into $\mu_{1}\,\mathcal{U}\,\square\mu_{2}$.)
Our method alternates parent selection and offspring generation, avoiding duplicates and enforcing temporal consistency by making every non-root node to have an ancestor with a temporal operator. For example, we disallow $\mu_{1}\land\square\mu_{2}$, since $\mu_{1}$ lacks a temporal operator ancestor. We simplify logic trees using rules such as $\square\square=\square$ to encourage logical simplicity. Mining stops when the current logic accepts all expert demonstrations and rejects all other trajectories. In practice, we limit the number of mining steps to $4$ for the wiping task and $5$ for the other tasks.
@@ IV-C Policy player: logic-constrained reinforcement learning
We introduce logic-constrained RL (Logic-CRL), which trains a policy to satisfy the TLTL constraints discovered by GA-TL-Mining. A main challenge is that TLTL constraints are non-Markovian; Their truths depend on entire state sequences rather than individual states. To resolve the mismatch, we embed each constraint into a product CMDP (PCMDP) by augmenting states with the TLTL constraint-based automaton states, thereby restoring the Markov property. Another challenge is the non-differentiable nature of logic constraints. To figure it out, we adopt Lagrangian CRL [tessler2018reward] with a novel state-wise constraint redistribution scheme, enabling conventional CRL frameworks to simplify the non-differentiable problem while still satisfying the non-Markovian TLTL constraint. We detail both processes in the following.
PCMDP: We construct a PCMDP by synchronously combining the CMDP’s transition system with a deterministic finite automata (DFA) converted from a TLTL constraint $\phi$. The DFA is the tuple $(\mathcal{Q}_{\phi},\Sigma_{\phi},\delta_{\phi},q_{0,\phi},F_{\phi})$, where $\mathcal{Q}_{\phi}$ is the set of states, $\Sigma_{\phi}=2^{\textbf{AP}}$ is the power set of APs, $\delta_{\phi}:\mathcal{Q}_{\phi}\times\Sigma_{\phi}\rightarrow\mathcal{Q}_{\phi}$ is the deterministic transition function, $q_{0,\phi}\in\mathcal{Q}_{\phi}$ is the initial state, and $F_{\phi}\subset\mathcal{Q}_{\phi}$ is the set of accepting states.
@@
Finally, the PCMDP formulation synthesizes a policy that maximizes cumulative rewards subject to the TLTL constraint $\phi$. However, optimizing policies with TLTL constraints presents challenges due to their sparse nature and the delayed evaluation at the end of episodes.
Lagrangian CRL: We introduce a Lagrangian CRL to handle sparse and temporally delayed TLTL constraint evaluation in policy learning. The sparse nature comes from Boolean semantics of TLTLs. To figure it out, we incorporate quantitative semantics with robustness function $\rho(\xi,\phi)$ to define a bounded but dense constraint-cost function $C_{\phi}^{\prime}\in[0,1]$, updating the discrete function $C_{\phi}$ in Eq. (5), as:
@@
where $\alpha\in[0,1]$ is a mixing weight, $\Upsilon\in\mathbb{R}^{+}$ limits the impact of large constraint violations, and $\mathrm{CLIP}$ truncates the negative robustness value within the range $0,\Upsilon]$.
To address the delayed evaluation, we introduce a redistribution scheme that spreads out the trajectory-level constraint cost $C^{\prime}_{\phi}(\xi)$ uniformly across all state-action pairs $({\bf s}_{t},{\bf a}_{t})\in\xi$, analogous to the uniform reward redistribution in IRCR [gangwani2020learning]. The resulting stepwise constraint cost $C^{\prime}_{\phi}({\bf s}_{t},{\bf q}_{t},{\bf a}_{t})$ is in the augmented state space $({\bf s}_{t},{\bf q}_{t})\in\mathcal{S}_{\phi}$ as
$C^{\prime}_{\phi}({\bf s}_{t},{\bf q}_{t},{\bf a}_{t})=\mathbb{E}_{\xi\sim\mathcal{D}}[C^{\prime}_{\phi}(\xi)|(({\bf s}_{t},{\bf q}_{t}),{\bf a}_{t})\in\xi],$ (11)
where $\mathcal{D}$ is the replay buffer containing previously collected trajectories in the augmented state space. This trajectory-space smoothing converts the trajectory constraint in Eq. (5) into a state-action constraint, allowing its use in Lagrangian CRL.
We formulate the Lagrangian dual of Eq. (5) to solve for the constrained-optimal $\pi_{k}$ in the $k$-th Logic-CRL process:
@@
where $\lambda$ is a learnable Lagrangian multiplier and $d$ is a constraint-cost threshold. We define a large threshold, $d=\frac{\varepsilon\alpha}{N_{\xi}}$, for stable optimization while enforcing the satisfaction of the binary logic constraint $C_{\phi}$, where $\varepsilon<1$ is a positive constant and $N_{\xi}$ is the number of trajectories sampled from the current policy $\pi$. To solve Eq. (12), we adopt SAC-Lag [ha2021learning]—a soft actor-critic (SAC) [haarnoja2018soft] version of a Lagrangian CRL method.
Further, to enhance off-policy learning in Eq. (12), we initialize the replay buffer $\mathcal{D}$ with the expert trajectories $\Xi_{E}$ and the previously sampled trajectories $\Xi_{1:k-1}$ that already satisfy the given constraint. After obtaining $\pi_{k}$, we sample zero-violation trajectories $\Xi_{k}$ and update $\Xi_{1:k}=\Xi_{1:k-1}\cup\Xi_{k}$, which serves as input to the next GA-TL-Mining step.
Lastly, after $N$ alternations of GA-TL-Mining and Logic-CRL, ILCL returns the best constraint. The two-player game progressively finds constraints toward the ground truth when the constraint space is convex, and the constraint player’s loss is linear and regularized [kim2024learning]. Although the logic constraint space is non-convex, we regularize the constraint player with tree size penalty as Eq. (6) to stabilize learning. We further choose the constraint $C_{\phi,i}$ that yields the most expert-like behavior from the obtained constraint set $\{C_{\phi,0},\dots,C_{\phi,N}\}$. For each candidate constraint, we evaluate its associated policy $\pi_{i+1}$ on the expert trajectories $\Xi_{E}$ and select the policy that minimizes the error to the expert actions,
$\pi_{i+1}=\underset{\pi_{i+1}\in[\pi_{1},\ldots,\pi_{N+1}]}{\text{argmin}}\mathbb{E}_{\begin{subarray}{c}(s_{E},a_{E})\sim\Xi_{E}\\ a\sim\pi_{i+1}(\cdot\mid s_{E})\end{subarray}}\,\|a_{E}-a\|_{2},$ (13)
and report the corresponding $C_{\phi,i}$.
@@ V Experimental Setup
We conduct simulated benchmarks and demonstrate real-world transfer of the learned constraints.
@@ V-A Benchmark Setups
1) Navigation: This task requires a point agent at $\mathbf{x}^{Agt}\in\mathbb{R}^{2}$ navigating to a goal position $\mathbf{x}^{Goal}\in\mathbb{R}^{2}$ while satisfying temporal constraints related to colored regions randomly placed in the workspace (see Fig. 4 Left). Let $\mathbf{p}^{R}$, $\mathbf{p}^{G}$, and $\mathbf{p}^{B}$ denote the polar coordinates $(p_{dist},p_{ang})\in\mathbb{R}^{2}$ of the nearest red, green, and blue regions, respectively, measured relative to the agent’s coordinate. For example, ${\bf p}^{R}=(p^{R}_{dist},p^{R}_{ang})$. We model this environment as two CMDPs, each subject to a distinct constraint, $\phi_{\text{GT1}}$ or $\phi_{\text{GT2}}$:
@@
2) Wiping: This task requires a roller agent at ${\bf x}^{Agt}\in\mathbb{R}^{2}$ wiping the ground surface between the start point $A$ (green) and the goal point $B$ (red), as illustrated in Fig. 4 Middle. Let $f_{c}$ be the contact force between the roller and the surface. $^{A}{\bf x}^{Agt}$ and $^{B}{\bf x}^{Agt}$ denote the positions of the agent relative to the start and goal points, respectively. For example, $^{A}{\bf x}^{Agt}=(^{A}x^{Agt},^{A}y^{Agt})$. We model this environment as a CMDP:
@@
  • $R=R_{goal}+R_{contact}+R_{penalty}$; $R_{goal}=\mbox{$\mathbf{1}$}(|^{B}x^{Agt}|<0.1)-|^{B}x^{Agt}|\cdot\mbox{$\mathbf{1}$}(|^{B}x^{Agt}|>0.1)$ is a goal reward, $R_{contact}=2\cdot\mbox{$\mathbf{1}$}(f_{c}>0.05)$ is a contact reward to encourage wiping contact, and $R_{penalty}=-0.2\cdot f_{c}^{2}$ is to penalize large contact.
@@
3) Peg-in-shallow-hole: A gripper agent requires to insert a thin peg into a shallow-angled hole while maintaining side contact (see Fig. 4 Right). Let $\Sigma_{h}$ be the 2D hole frame tilted by angle $\theta^{h}$. The gripper at $^{\Sigma_{h}}{\bf x}^{grip}$ is $(^{\Sigma_{h}}x^{grip},^{\Sigma_{h}}y^{grip},^{\Sigma_{h}}\theta^{grip})\in\mathbb{R}^{3}$, where $\theta^{grip}$ is the gripper tilt in the hole frame. The peg is at $^{\Sigma_{h}}{\bf x}^{peg}$. We also define two features $d_{hole}^{peg}$ and $d_{jaw}^{peg}$ as the peg edge-to-hole and peg-edge-to-support-jaw distances, respectively. Then, a CMDP is
@@
  • $R=(1-\tanh(5|^{\Sigma_{h}}y^{peg}-^{\Sigma_{h}}y^{peg*}|))+5\cdot\mbox{$\mathbf{1}$}(^{\Sigma_{h}}y^{peg}-^{\Sigma_{h}}y^{peg*})$, where $^{\Sigma_{h}}y^{peg}$ and $^{\Sigma_{h}}y^{peg*}$ are the current and desired vertical position of the peg in terms of the hole frame $\Sigma_{h}$.